# Dissecting CLIP: Decomposition with a Schur Complement-based Approach

### v2

@@ Appendix D Results with Naive Text Embedding Subtraction without considering the adjust matrix $\Gamma^{*}$
Figure 23: Effect of removing encoded "cassette player" text on top of ImageNet samples. Top figure represents naive method and bottom figure represents SC method.
@@
Figure 25: Kernel PCA clusters before and after CLIP correction on the captioned ImageNet dataset, comparing the SCE decomposition approach with the naive approach.
Appendix E Additional Numerical Results
We evaluated several text-to-image models using the SCE metric to assess their performance. Figure 26 summarizes our findings for DALL-E 2 [32], DALL-E 3 [7], Kandinsky 3 [33], and FLUX.1-schnell [27], tested on 5,000 MSCOCO [34] captions.
@@ Appendix E Additional Numerical Results
Figure 26: Comparison of different text-to-image models with Vendi-1.0 (generated image diversity) and SCE (image generator diversity).

### v3

@@ Appendix G Results with Naive Text Embedding Subtraction without considering the adjustment matrix $\Gamma^{*}$
Figure 24: Effect of removing encoded ”cassette player” text on top of ImageNet samples. Top figure represents naive method and bottom figure represents SC method.
@@
Figure 26: Kernel PCA clusters before and after CLIP correction on the captioned ImageNet dataset, comparing the Scendi decomposition approach with the naive approach.
Appendix H Additional Numerical Results
We evaluated several text-to-image models using the Scendi metric to assess their performance. Figure 27 summarizes our findings for DALL-E 2 [47], DALL-E 3 [40], Kandinsky 3 [48], and FLUX.1-schnell [29], tested on 5,000 MSCOCO [35] captions.
@@ Appendix H Additional Numerical Results
Figure 27: Comparison of different text-to-image models with Vendi-1.0 (generated image diversity) and Scendi (image generator diversity).
Figure 28: t-SNE plot of animals with objects dataset.